OpenStudy (anonymous):

Event B is dependent on event A, and event A occurs before event B. Which formula can be used to find the probability of event A?

OpenStudy (kirbykirby):

You can use a conditional probability: \[P(B|A)=\frac{P(B \cap A)}{P(A)}\] You know know A occured first, so you are given info about A. That is where P(B|A) comes from. Then you can easily solve for P(A) by isolating it.

OpenStudy (anonymous):

How to isolate?

OpenStudy (kirbykirby):

@kevinhunterwood69 Well, let's think of this as a fractional expression that maybe looks simpler in which we want to isolate \(y\), like: \[ x=\frac{4}{y}\], , this is equivalent to saying \[ \frac{x}{1}=\frac{4}{y}\] What you can do is cross-multiply to get: \[ xy=4\cdot 1\] To isolate \(y\), just divide both sides by \(x\) :) Now try the same procedure with the conditional probability formula, but just replace \(x\) with \(P(B|A)\), \(4\) with \(P(B \cap A)\) and \(y\) with \(P(A)\).

OpenStudy (anonymous):

Ah okay thank you @kirbykirby

OpenStudy (anonymous):

wait whats the answer lol